# Robust Multimodal Learning with Missing Modalities via Parameter-Efficient Adaptation

### v3

@@ Abstract
Multimodal learning seeks to utilize data from multiple sources to improve the overall performance of downstream tasks. It is desirable for redundancies in the data to make multimodal systems robust to missing or corrupted observations in some correlated modalities. However, we observe that the performance of several existing multimodal networks significantly deteriorates if one or multiple modalities are absent at test time. To enable robustness to missing modalities, we propose a simple and parameter-efficient adaptation procedure for pretrained multimodal networks. In particular, we exploit modulation of intermediate features to compensate for the missing modalities. We demonstrate that such adaptation can partially bridge performance drop due to missing modalities and outperform independent, dedicated networks trained for the available modality combinations in some cases. The proposed adaptation requires extremely small number of parameters (e.g., fewer than 0.7% of the total parameters) and applicable to a wide range of modality combinations and tasks. We conduct a series of experiments to highlight the missing modality robustness of our proposed method on 5 different datasets for multimodal semantic segmentation, multimodal material segmentation, and multimodal sentiment analysis tasks. Our proposed method demonstrates versatility across various tasks and datasets, and outperforms existing methods for robust multimodal learning with missing modalities.
@@ 1 Introduction
In this paper, we propose a parameter-efficient approach to adapt existing multimodal networks to perform well on different missing modality scenarios. Our main objective is to modify the network in a controllable manner as a function of available modalities. For instance, if a modality is missing, we seek to modify how the features from available modalities are extracted and fused for the inference. Instead of learning an independent network for each modality combination, our goal is to perform parameter-efficient adaptation. Figure 1 illustrates our proposed method, where a given multimodal network can be adapted to arbitrary modality combinations by transforming the intermediate features of the available input modalities at different layers. To achieve parameter-efficient adaptation, we propose to use simple linear transformations such as scaling and shifting to modulate the intermediate features or low-rank increments of features to compensate for the missing modalities. Our method does not require retraining the entire model or any specialized training strategy. The adapted networks provide significant performance improvement over the multimodal networks trained with all modalities and tested with missing modalities. Performance of the adapted models is also comparable or better than the models that are exclusively trained for each input modality combination. We present a series of experiments to evaluate our method and compare with existing methods for robust MML. We tested different parameter-efficient adaptation strategies and found intermediate feature modulation with scaling and shifting provides overall best performance with less than 0.7% of the total parameters.
@@
  • The adapted networks provide notably improved performance with missing modalities when compared to models trained with all modalities and is comparable to or better than the networks trained for specific modality combinations.
  • Our approach is versatile and adaptable to a wide range of multimodal tasks and models. Detailed evaluations on different datasets and tasks show that our approach outperforms existing baseline methods and robust models designed for specific tasks and datasets.
Figure 1: a) Overview of our approach for robust multimodal learning with missing modalities via parameter-efficient adaptation. A model pretrained on all the modalities is adapted using a small number of learnable parameters to handle different modality combinations. We insert adaptable layers after each layer of the encoders and the fusion block to learn the modulation as a function of the available input modalities to compensate for the missing modalities. The grayed-out branch (missing modality branch) is inactive and does not contribute to the output. b) Low-rank model adaption computes features using frozen weights and low-rank weight updates and combine them. c) Scale and shift feature adaptation transforms input by element-wise multiplication and addition. One set of parameters is learned for each modality combination.
@@ 2 Related Work
Knowledge distillation and generation methods have also become popular for robust MML. Studies by [47] and [64] used GAN based generative models while [11] used VAE based generative models for imputing missing modalities from available input modalities for unlerlying multimodal tasks. Recently [44] introduced an approach to learn missing modality tokens from available modalities. Different knowledge distillation approaches have also been applied in several multimodal tasks. [52] proposed mean teacher and [40] introduced multimodal teacher for semi-supervised image segmentation. [48] and [18] applied self-distillation loss for robust RGB-thermal semantic segmentation. Apart from these approaches, weight space ensembling [59], policy learning [38] and optimal fusion strategy designing [40] were also studied for robust MML for various tasks.
@@
Parameter-efficient network adaptation has become very popular for uni-modal network adaptation in recent years [23, 34]. A number of parameter-efficient methods have been proposed for transfer learning [20, 10] and domain/task adaptation [24, 31]. We can divide the approaches into two major categories: low-rank/additive approach and feature transformation approach.
Low-rank/additive adaptation has been applied for uni-modal model fine-tuning and domain adaptation. For instance, LoRA [24], QLoRA [9], KronA [12] and KAdaptataion [22] learn low-rank factors for task/domain adaptation. Let $W$ be the weight matrix of any dense layer of a given pretrained uni-modal model. These approaches learn a low-rank weight update matrix $\Delta W$ to transform the input $x$ to that layer as $h=Wx+\Delta Wx$, where $h$ is the updated feature. Since the update matrix $\Delta W$ is low-rank, the number of learnable parameters remains a fraction of the total number of model parameters. BitFit [5] is another approach that adapts the bias terms while keeping everything else frozen.
@@ 3.1.2 Parameter-efficient approach
The adaptation method shown in (3) is considered parameter-efficient if the number of parameters in $\Delta_{\mathcal{S}}$ is significantly smaller compared to the total number of parameters in $\Theta_{\mathcal{M}}$. During adaptation, we only train the learnable parameters $\Delta_{\mathcal{S}}$. In our experiments, as we discuss later, we keep $\Theta_{\mathcal{M}}$ frozen and demonstrate that less than $0.7\%$ of the total parameters for $\Delta_{\mathcal{S}}$ are sufficient for network adaptation (see supplementary material for details).
@@ 3.1.3 Need for parameter-efficient adaptation
In the last few years, a number of approaches have been proposed for MML with missing modalities. To the best of our knowledge, parameter-efficient adaptation is still unexplored in this field. The current methods for robust MML, as discussed in Section 2, require retraining the whole model with specialized training strategy [48, 25] or utilize extra module/sub-network to guide the multi-modal model [40, 52]. Furthermore, these methods are not very generic and do not perform well on different missing modality scenarios as shown in Table 2 and 3. To solve these issues, we propose parameter-efficient adaptation for enhancing robustness of MML in the face of missing modalities. Our approach requires learning a very small number of parameters for different missing modality scenarios without the need to retrain the whole network. Furthermore, it is also applicable to diverse model architectures, tasks and modality combinations.
3.2 Parameter-Efficient Adaptation for Missing Modalities
In this section we first discuss how we adapt a given multimodal network for different missing modality scenarios. Then we discuss the reasons behind selecting intermediate feature modulation as the adaptation technique. Finally, we compare it with other parameter-efficient adaptation methods and discuss about some key benefits of our selected approach.
Adaptation for multimodal models. To the best of our knowledge, no parameter-efficient adaptation approach has been proposed or applied for multimodal model adaptation to handle missing modalities. We draw our motivation from low-rank adaptation and feature modulation based approach to compensate for the missing modalities. For instance, LoRA [24], BitFit [5], layer/feature normalization [1, 26], and scale and shift features (SSF) [31] approaches can enhance the representation capabilities of deep models for domain adaptation and model fine-tuning. We extend these adaptation approaches to build a generic framework that can transform the intermediate features of the available modalities to find an optimal feature representation to compensate for the performance gap due to missing modalities.
@@ 3.2.1 Training: model adaptation for missing modalities
Our approach for multi-modal model adaptation for missing modality is illustrated in Figure 1. Without loss of generality, let us assume a generic multi-modal model in which each modality goes through a separate modality-specific encoder. The extracted features from all the modalities are fused with a fusion block and the fused feature is passed to a decoder head for making prediction. Each encoder and the fusion block have multiple layers (linear/convolutional/norm layers). This setup can be easily generalized to models with shared encoder and/or different (early or mid) fusion strategy. We also do not make any assumption about the underlying model architecture.
@@
Below we discuss how to use low-rank and intermediate feature modulation-based multimodal network adaptation to accommodate missing modalities. We believe our framework is generic and can also incorporate other parameter-efficient adaptation approaches.
@@
where $h_{m,o}$ is the transformed output feature that is passed to the next layer in the model. Since the weight update matrix $\Delta W_{m}$ is low-rank, the total number of learnable parameters remains a fraction of the total number of model parameters. We can represent the learnable parameters $\Delta_{\mathcal{S}}=\{\Delta W_{m}\}_{m\in\mathcal{S}}$ as the collection of all low-rank update matrices.
Intermediate feature modulation. We extend SSF [31] method to work with multimodal models with missing modalities. The adaptable SSF layers modulate the intermediate tokens/features from each available modality at every layer as shown in Figure 1c. For the $m^{th}$ input modality where $m\in\mathcal{S}$, we denote the learnable scale parameters as $\gamma_{m}\in\mathbb{R}^{d}$ and the learnable shift parameters as $\beta_{m}\in\mathbb{R}^{d}$ where $d$ is the embedding dimension of the model. Since we insert SSF layers after each frozen layer in the pretrained model $f$, the output ${h}_{m,o}\in\mathbb{R}^{N\times d}$ form any frozen layer for the $m^{th}$ input modality goes through the SSF layer that follows it. The SSF layer applies a linear transformation on ${h}_{m,o}$. We can formulate the linear transformation for multimodal input as follows:
@@
where ${h}_{m,i}\in\mathbb{R}^{N\times d}$ is the transformed feature which is fed to the next frozen layer in the model and $N$ is the number of tokens. Note that if the output of any layer is of shape $(H\times W\times d)$ (for convolutional layers), we reshape it to $(N\times d)$, where $N=H\times W$, before applying the linear transformation. We reshape the transformed feature back to the original shape (if required) before passing it to the next layer. We can represent the learnable parameters as $\Delta_{\mathcal{S}}=\{\gamma_{\mathcal{S}},\beta_{\mathcal{S}}\}=\{\gamma_{m},\beta_{m}\}_{m\in\mathcal{S}}$. BitFit [5] method can also be used for adaptation as we only need to learn the bias/shift terms $\beta_{m}$ for all $m\in\mathcal{S}$. We modify (5) as
@@ 3.2.2 Inference: model adaptation for missing modalities
Since we are inserting the adaptable layers after each layer, it does not require any major change to the model architecture and can be done easily without reloading all the model parameters $\Theta_{\mathcal{M}}$. We just need to load the parameters in $\Delta_{\mathcal{S}}$ and insert them into the model. Since we only insert a very small number of additional parameters (less than 0.7% of the total model parameters), it adds very limited computational overhead. Furthermore, if a different subset of modalities becomes available, the adjustment is straightforward. We only need to replace the existing learned parameters $\Delta_{\mathcal{S}}$ with the corresponding parameters for the available modality set, ensuring adaptability and flexibility in handling diverse combinations of available modalities during the testing phase.
@@ 3.2.3 Feature modulation vs low-rank adaptation
While we present three adaptation approaches in (4), (5), and (6), we select intermediate feature modulation with SSF (5) as the main approach for our experiments. We primarily selected this technique for robust MML with missing modalities because of its simplicity and effectiveness. Our experiments show that feature transformation via simple linear transformation with SSF works well for most of the scenarios compared to other parameter-efficient adaptation approaches as summarized in Table 4. We provided a detailed comparison in terms of mean accuracy, F1 score and % mIoU in Table 8, 9 and 10 in the supplementary section. SSF shows great promise in enhancing representation power [1], faster convergence [26], prevents loss of information in the representation learning process [60] and mitigates distribution mismatch between the upstream and downstream tasks [31]. These characteristics motivated us to extend this method for multimodal learning with missing modalities and build a generic framework that is very effective in learning the proper modulation of available input modalities to bridge the performance gap in the face of missing modalities.
@@ 4 Experiments and Results
We performed detailed experiments to evaluate the performance of our proposed method for multimodal semantic segmentation, multimodal material segmentation and multimodal sentiment analysis tasks on five datasets. We also present comparison with existing baseline methods that are robust to missing modalities.
Table 1: Performance of Pretrained, Dedicated, and Adapted networks with missing modalities. CMNeXt is the base model for multimodal semantic segmentation for MFNet and NYUDv2 datasets and multimodal material segmentation for MCubeS dataset. HHA-encoded images were used instead of raw depth maps. Bold letters represent best results.
Dataset Input Missing Pretrained Dedicated Adapted (Ours)
MFNet RGB-Thermal - 60.10 60.10 -
RGB Thermal 53.71 55.86 55.22
Thermal RGB 35.48 53.34 50.89
NYUDv2 RGB-Depth - 56.30 56.30 -
RGB Depth 51.19 52.18 52.82
Depth RGB 5.26 33.49 36.72
MCubeS RGB-AoLP-DoLP-NIR - 51.54 51.54 -
RGB-AoLP-DoLP NIR 49.06 49.48 51.11
RGB-AoLP DoLP-NIR 48.81 48.42 50.62
RGB AoLP-DoLP-NIR 42.32 48.16 50.43
4.1 Datasets
Multimodal semantic segmentation. We experimented on two datasets for multimodal semantic segmentation tasks: MFNet dataset [17] with RGB-Thermal input modalities and NYUDv2 dataset [49] with RGB-Depth input modalities. These datasets are divided into train and test sets along with ground truth per-pixel annotation for the underlying semantic segmentation tasks.
Multimodal material segmentation. We used MCubeS dataset [32] for multimodal material segmentation having RGB, Angle of Linear Polarization (AoLP), Degree of Linear Polarization (DoLP) and Near-Infrared (NIR) images as input modalities. The dataset is divided into train, validation and test sets along with ground truth per-pixel annotation for 20 material classes.
Multimodal sentiment analysis. CMU-MOSI from [69] and CMU-MOSEI from [3] are two popular datasets for multimodal sentiment analysis using audio, visual and text as input modalities. They contain 2199 and 23453 annotated data samples, respectively, divided into train, validation, and test sets.
Figure 2: Examples of predicted segmentation maps for the Pretrained and Adapted models for multimodal semantic segmentation (on MFNet and NYUDv2) and material segmentation (on MCubeS). Title above each subimage shows method name (tested with available modalities). CMNeXt column shows the predictions with all the modalities. Segmentation quality improves significantly after model adaptation for all the input modality combinations. Green boxes highlight areas with salient differences in results (e.g., cars and humans missing in the Pretrained model with missing modalities but visible in the Adapted model). For MCubeS dataset, we only show RGB images and A, D and N denote angle of linear polarization, degree of linear polarization, and near-infrared, respectively.

### v4

@@ Abstract
Multimodal learning seeks to utilize data from multiple sources to improve the overall performance of downstream tasks. It is desirable for redundancies in the data to make multimodal systems robust to missing or corrupted observations in some correlated modalities. However, we observe that the performance of several existing multimodal networks significantly deteriorates if one or multiple modalities are absent at test time. To enable robustness to missing modalities, we propose a simple and parameter-efficient adaptation procedure for pretrained multimodal networks. In particular, we exploit modulation of intermediate features to compensate for the missing modalities. We demonstrate that such adaptation can partially bridge performance drop due to missing modalities and outperform independent, dedicated networks trained for the available modality combinations in some cases. The proposed adaptation requires extremely small number of parameters (e.g., fewer than 1% of the total parameters) and applicable to a wide range of modality combinations and tasks. We conduct a series of experiments to highlight the missing modality robustness of our proposed method on five different multimodal tasks across seven datasets. Our proposed method demonstrates versatility across various tasks and datasets, and outperforms existing methods for robust multimodal learning with missing modalities.
@@ 1 Introduction
Figure 1: a) Overview of our model adaptation approach for robust MML. A model pretrained on all the modalities is adapted using a small number of learnable parameters to handle different modality combinations. We insert adaptable layers after each layer of the encoders and the fusion block to learn the modulation as a function of the available input modalities to compensate for the missing modalities. The grayed-out branch (missing modality) is inactive and does not contribute to the output. b) Low-rank model adaption computes features using frozen weights and low-rank weight updates and combine them. c) Scale and shift feature adaptation transforms input by element-wise multiplication and addition.
In this paper, we propose a parameter-efficient approach to adapt existing multimodal networks to perform well on different missing modality scenarios. Our main objective is to modify the network in a controllable manner as a function of available modalities. For instance, if a modality is missing, we seek to modify how the features from available modalities are extracted and fused for the inference. Instead of learning an independent network for each modality combination, our goal is to perform parameter-efficient adaptation. Figure 1 illustrates our proposed method, where a given multimodal network can be adapted to arbitrary modality combinations by transforming the intermediate features of the available input modalities at different layers. To achieve parameter-efficient adaptation, we propose to use simple linear transformations such as scaling and shifting to modulate the intermediate features or low-rank increments of features to compensate for the missing modalities. Our method does not require retraining the entire model or any specialized training strategy. The adapted networks provide significant performance improvement over the multimodal networks trained with all modalities and tested with missing modalities. Performance of the adapted models is also comparable or better than the models that are exclusively trained for each input modality combination as shown in Table 1. We present a series of experiments to evaluate our method and compare with existing methods for robust MMLon five multimodal tasks across seven datasets (Section 4.1). We tested different parameter-efficient adaptation strategies and found intermediate feature modulation with scaling and shifting provides overall best performance, which is discussed in Section S4. Our method shows significant performance improvement with less than 1% additional learnable parameters as discussed in Section 4.7.2.
@@
  • The adapted networks provide notably improved performance with missing modalities when compared to models trained with all modalities and is comparable to or better than the networks trained for specific modality combinations (Table 1).
  • Our approach is versatile and adaptable to a wide range of multimodal tasks, datasets and models. Detailed evaluations on different datasets and tasks show that our approach outperforms existing baseline methods and robust models designed for specific tasks and datasets (Section 4.3 - Section 4.6).
@@ 2 Related Work
Knowledge distillation and generation methods have also become popular for robust MML. Studies by [25] and [24] used GAN based generative models while [26] used VAE based generative models for imputing missing modalities from available input modalities for unlerlying multimodal tasks. Recently [36] introduced an approach to learn missing modality tokens from available modalities. Different knowledge distillation approaches have also been applied in several multimodal tasks. [21] proposed mean teacher and [22] introduced multimodal teacher for semi-supervised image segmentation. [20] and [15] applied self-distillation loss for robust RGB-thermal semantic segmentation. Apart from these approaches, weight space ensembling [37], policy learning [14], optimal transport based approach [38] and optimal fusion strategy designing [22] were also studied for robust MML for various tasks.
@@
Parameter-efficient network adaptation has become very popular in recent years [39, 40]. A number of parameter-efficient methods have been proposed for transfer learning [41, 42] and uni-modal domain/task adaptation [43, 44]. We can divide the approaches into following two major categories:
Low-rank/additive adaptation has been applied for uni-modal model fine-tuning and domain adaptation. For instance, LoRA [43], QLoRA [45], KronA [46] and KAdaptataion [47] learn low-rank factors for task/domain adaptation. Let $W$ be the weight matrix of any dense layer of a given pretrained uni-modal model. These approaches learn a low-rank weight update matrix $\Delta W$ to transform the input $x$ to that layer as $h=Wx+\Delta Wx$, where $h$ is the updated feature. Since the update matrix $\Delta W$ is low-rank, the number of learnable parameters remains a fraction of the total number of model parameters.
@@ 3.1.2 Parameter-efficient approach
The adaptation method shown in (3) is considered parameter-efficient if the number of parameters in $\Delta_{\mathcal{S}}$ is significantly smaller compared to the total number of parameters in $\Theta_{\mathcal{M}}$. During adaptation, we keep $\Theta_{\mathcal{M}}$ frozen and demonstrate that less than $1\%$ of the total parameters for $\Delta_{\mathcal{S}}$ are sufficient for network adaptation (Section 4.7.2).
@@ 3.1.3 Need for parameter-efficient adaptation
In recent years, a number of approaches have been proposed for MML with missing modalities. To the best of our knowledge, parameter-efficient adaptation is still unexplored in this field. The current methods for robust MML, as discussed in Section 2, require retraining the whole model with specialized training strategy [20, 18] or utilize extra module/sub-network to guide the multi-modal model [22, 21]. Furthermore, these methods are not very generic and do not perform well on different missing modality scenarios as shown in Table 2 and 3. To solve these issues, we propose parameter-efficient adaptation for enhancing missing modality robustness of MML. Our approach requires learning a very small number of parameters for different missing modality scenarios without the need to retrain the whole network. Furthermore, it is also applicable to diverse model architectures, tasks and modality combinations as discussed in Section 4.
3.2 Parameter-Efficient Adaptation for Robust MML
This section outlines our approach for multimodal network adaptation for missing modalities. We explain the reasons behind selecting intermediate feature modulation and compare with other parameter-efficient methods, highlighting key benefits of our approach.
Adaptation for multimodal models. To the best of our knowledge, no parameter-efficient adaptation approach has been proposed or applied for multimodal model adaptation to handle missing modalities. We draw our motivation from low-rank adaptation [43, 51, 49, 48] and feature modulation based approach [44].
These approaches can enhance the representation capabilities of deep models. We extend these adaptation approaches to build a generic framework that can transform the intermediate features of the available modalities to find an optimal feature representation to compensate for the performance gap due to missing modalities.
@@ 3.2.1 Training: model adaptation for missing modalities
Our approach is illustrated in Figure 1. Without loss of generality, let us assume a generic multi-modal model in which each modality goes through a separate encoder for feature extraction, followed by a fusion block to fuse the extracted features. The fused feature is passed to a decoder head for making prediction. This setup can be easily generalized to models with shared encoder, different encoder/model architecture and/or different (early or mid) fusion strategy.
@@
Below we discuss how to use low-rank and intermediate feature modulation-based multimodal network adaptation to accommodate missing modalities. Our framework is generic and can also incorporate other parameter-efficient adaptation approaches.
@@
where $h_{m,o}$ is the transformed output feature that is passed to the next layer in the model. Since $\Delta W_{m}$ is low-rank, the total number of learnable parameters remains a fraction of the total number of model parameters. We can represent the learnable parameters $\Delta_{\mathcal{S}}=\{\Delta W_{m}\}_{m\in\mathcal{S}}$ as the collection of all low-rank update matrices.
Intermediate feature modulation. We extend SSF [44] method to work with multimodal models with missing modalities. The adaptable SSF layers modulate the intermediate tokens/features from each available modality at every layer as shown in Figure 1c. For the $m^{th}$ input modality where $m\in\mathcal{S}$, we denote the learnable scale and shift parameters as $\gamma_{m}\in\mathbb{R}^{d}$ and $\beta_{m}\in\mathbb{R}^{d}$ respectively where $d$ is the embedding dimension of the model. The output ${h}_{m,o}\in\mathbb{R}^{N\times d}$ form any frozen layer for the $m^{th}$ input modality goes through the SSF layer that follows it. The SSF layer applies a linear transformation on ${h}_{m,o}$ as follows:
@@
where ${h}_{m,i}\in\mathbb{R}^{N\times d}$ is the transformed feature which is fed to the next frozen layer in the model and $N$ is the number of tokens. Note that if the output of any layer is of shape $(H\times W\times d)$ (for convolutional layers), we reshape it to $(N\times d)$, where $N=H\times W$, before applying (5). We reshape the transformed feature back to the original shape (if required) before passing it to the next layer. We can represent the learnable parameters as $\Delta_{\mathcal{S}}=\{\gamma_{\mathcal{S}},\beta_{\mathcal{S}}\}=\{\gamma_{m},\beta_{m}\}_{m\in\mathcal{S}}$. BitFit [51] method can also be used for adaptation as we only need to learn the bias/shift terms $\beta_{m}$ for all $m\in\mathcal{S}$. We modify (5) as
@@ 3.2.2 Inference: model adaptation for missing modalities
Since we are inserting the adaptable layers after each layer, it does not require any major change to the model architecture and can be done easily without reloading all the model parameters $\Theta_{\mathcal{M}}$. We just need to load the parameters in $\Delta_{\mathcal{S}}$ and insert them into the model. Since we only insert a very small number of additional parameters, it adds very limited computational overhead. Furthermore, if a different subset of modalities becomes available, the adjustment is straightforward. We only need to replace the existing learned parameters $\Delta_{\mathcal{S}}$ with the corresponding parameters for the available modality set, ensuring adaptability and flexibility in handling diverse combinations of available modalities during the testing phase.
@@ 3.2.3 Feature modulation vs low-rank adaptation
While we present three adaptation approaches in (4), (5), and (6), we select intermediate feature modulation with SSF (5) as the main approach for our experiments. We primarily selected this technique because of its simplicity and effectiveness. Our experiments show that feature transformation via simple linear transformation with SSF works well for most of the scenarios compared to other parameter-efficient adaptation approaches as summarized in Table S3. We provided a detailed comparison in terms of mean accuracy, F1 score and % mIoU in Table S4, S5 and S6 in the supplementary section. SSF shows great promise in enhancing representation power [49], faster convergence [48], prevents loss of information in the representation learning process [50] and mitigates distribution mismatch between the upstream and downstream tasks [44]. These characteristics motivated us to extend this method for MML with missing modalities and build a generic framework that is very effective in learning the proper modulation of available input modalities to bridge the performance gap in the face of missing modalities.
@@ 4 Experiments and Results
We performed detailed experiments to evaluate the effectiveness and generalizability of our proposed method on five multimodal tasks across seven datasets. In this section, we present comparison with existing baseline methods that are robust to missing modalities.
4.1 Datasets and Tasks
In this section, we provide a brief description of each dataset. Please refer to Section S1 in the supplementary materials for comprehensive details on each dataset.
4.1.1 Multimodal semantic segmentation.
MFNet dataset [53] contains 1569 registered RGB-Thermal image pairs and divided into train and test sets. Each image is $640\times 480$ pixels, contains annotation for 9 classes.
NYUDv2 dataset [54] has 1449 pairs of aligned RGB-Depth image pairs. It is divided into train and test sets having 795 and 654 image pairs respectively. Each image is $640\times 480$ pixels and contains annotation for 40 classes. We used HHA encoded images [55] instead of raw depth maps for our experiments.
4.1.2 Multimodal material segmentation
MCubeS dataset [56] has 4 input modalities: RGB, Angle of Linear Polarization (AoLP), Degree of Linear Polarization (DoLP) and Near-Infrared (NIR). The dataset is divided into train, validation and test sets containing 302, 96 and 102 sets of images respectively along with ground truth per-pixel annotation for 20 material classes.
4.1.3 Multimodal action recognition
NTU RGB+D (NTU60) dataset [57] contains 56,880 video samples across 60 action classes. It contains RGB videos $(1920\times 1080)$, depth map sequences $(512\times 424)$, infrared (IR) videos $(512\times 424)$ and 3D skeletal data. We use RGB and depth data for our experiments and evaluate performance using cross subject protocol.
4.1.4 Multimodal sentiment analysis
CMU-MOSI dataset [58] contains audio, visual and text modality for multimodal sentiment analysis. The dataset has 2,199 samples divided into train, validation and test containing 1,284, 229 and 686 samples respectively.
CMU-MOSEI dataset [59] is another large scale dataset. It contains 23,453 samples having audio, visual and text. The dataset is divided into train, validation and test sets.
4.1.5 Multimodal classification
UPMC Food-101 dataset [60] is a popular multimodal classification dataset containing image and text as input modalities. The dataset contains 90,704 image-text pairs and 101 food categories.